OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF and has a ratio of similarity of 10:3. If side EF measures 30 mm, how long is side BC?

OpenStudy (anonymous):

Please, please, please, please,pleaseee help :)

OpenStudy (anonymous):

without a picture i will assume that the labelling is in alphabetical order so \(\frac{BC}{EF}=\frac{10}{3}\)

OpenStudy (anonymous):

we can replace \(EF\) by \(30\) and get \[\frac{BC}{30}=\frac{10}{3}\] and therefore \[BC=\frac{30\times 10}{3}=10\times 10=100\]

OpenStudy (anonymous):

Oh, okayy. :D So the answer is 100mm?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank youu! Could you help me with one more thing?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Nvmind. I understand this one. Thank you so much, really♥

OpenStudy (anonymous):

yw